OpenStudy (anonymous):

Given the one-to-one function f(x)=(6x+1)/(x-2) , find f^-1(x) .

OpenStudy (anonymous):

set \[x=\frac{6y+1}{y-2}\] and solve for y

OpenStudy (anonymous):

i did that so many times i keep getting thesame answer but its wrong

OpenStudy (anonymous):

\[x(y-2)=6y+1\] \[xy-2x=6y+1\] \[xy-6y=1+2x\] \[(x-6)y=1+2x\] \[y=\frac{1+2x}{x-6}\] so \[f^{-1}(x)=\frac{1+2x}{x-6}\] check my algebra because i make mistakes

OpenStudy (anonymous):

okay one second thank you

OpenStudy (anonymous):

its right! thanks so much. I tried it so many times